OpenStudy (anonymous):

Equation for the tangent plane of z = sinx+cosy at (0,0)

OpenStudy (abb0t):

Do you know the general formula? I believe there is one.

OpenStudy (anonymous):

fx (x-x0) + fy (y-y0) or something along those lines I ended up getting cosx =0 which just looks wrong to me

OpenStudy (abb0t):

\(z-f(x_0,x_0)=f_x(x_0,y_0)(x-x_0)+f_y(x_0,x_0)(y-y_0)\)

OpenStudy (abb0t):

cos(0) = 1

OpenStudy (kc_kennylau):

\[\begin{array}{rl}f(x,y)&=&\sin x+\cos y\\f_x(x,y)&=&\cos x\\f_y(x,y)&=&-\sin y\end{array}\] \[\begin{array}{rl}\large f_x(x_0,y_0)(x−x_0)+f_y(x_0,y_0)(y−y_0)−(z−z_0)&=&0\\\cos0\cdot x-\sin 0\cdot y-[z-(\sin0-\cos0)]&=&0\\x-(z+1)&=&0\\x-z-1&=&0\end{array}\]

OpenStudy (kc_kennylau):

P.S. It's my first day doing this so I'm not sure if I'm right or not. (actually I just searched about this in a few moments ago)

OpenStudy (anonymous):

Let me check over

OpenStudy (abb0t):

It's pretty straight forward, just find the partial with respect to x, and y. Then plug in the given conditions and you get an eqaution :)

OpenStudy (kc_kennylau):

sorry it should be \(\sin0+\cos0\) which gives, at the end, \(x-z+1=0\)

OpenStudy (anonymous):

right. I think I've got it now. I'd give the medal to both of you if I could

OpenStudy (anonymous):

thank you both

OpenStudy (kc_kennylau):

I've given @abb0t a medal on behalf of you @rymdenbarn :)

OpenStudy (anonymous):

Thanks :p

OpenStudy (kc_kennylau):

no problem xD